A solution is made by mixing 244.0 mL 244.0 mL of ethanol initially at 13.0 ∘ C 13.0 ∘C with 244.0 mL 244.0 mL of water initially at 21.4 ∘ C 21.4 ∘C . What is the final temperature of the solution assuming that no heat is lost? The density of ethanol is 0.789 g/mL 0.789 g/mL and the density of water is 1.00 g/mL 1.00 g/mL . The specific heat of ethanol is 2.46 J/g·°C 2.46 J/g·°C and the specific heat of water is 4.184 J/g·°C 4.184 J/g·°C .
Answer #1

Answer:-

This question is answered by using the simple concept of calorimetry which gives that heat lost is equal to the heat gained.

The answer is given in the image,

Answer: V = 944.0m2 MI= 244 x 0.789 - 192.5169. V₂ = 244.0m m =244-0 X1.00 = 2449, C = 2.467/9.°C ₂ =4.184.5/9.06 Ty = 13.0°